# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v1

@@ Abstract
One of the most critical challenges for deploying distributed learning, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile devices. While it is a common belief that the major energy consumption of mobile devices comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile devices in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile devices by adaptively adjusting the number of decoding iterations. We theoretically prove that FL with communication errors can converge at the same rate as error-free communication as long as the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
@@ I-A Motivation
Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [2]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [3] to enable the training of deep learning models without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7]. Another issue for wireless FL is the energy consumption of mobile devices due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
@@
Many innovative approaches have been proposed to reduce the energy consumption of mobile devices in FL from different perspectives. For instance, reinforcement learning was utilized in [9] to identify a near-optimal clustering of participating devices to reduce the overall energy consumption of mobile devices. In [10], the trade-off between learning accuracy and users’ total energy consumption was optimized through computational and communication resource allocation. Besides system-level resource management, some works focused on the power consumption of specific hardware components such as transmitters, receivers, and artificial intelligence (AI) chips. Techniques such as model sparsification [11] and quantization [12, 13, 14, 15, 16, 17] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters. The authors of [18] and [19] minimized the transmission energy while maintaining the learning performance. From computation perspective, model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [20, 21, 22]. Additionally, some works, such as [10], [23], and [24], considered the joint optimization of the uplink transmission and local training energy consumption. However, very few works have paid attention to the energy consumption of channel decoder. As will be shown in Section II-B, channel decoders consume a substantial portion of energy because of the iterative process to decode massive DNN parameters.
@@
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in mobile clients and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient channel decoding strategy is designed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[25]. Furthermore, we conduct a theoretical analysis on the relation between the bit error rate (BER) and the convergence rate of FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption.
@@ I-B Contributions
This paper investigates the impact of channel decoding on the energy consumption of wireless FL. The major contributions are summarized as follows:
  • We theoretically analyze the relation between BER and the convergence rate of FL. The maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
  • Based on the analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to reduce the energy consumption of wireless FL while satisfying the BER requirement for a target convergence rate.
  • Experimental results show that the proposed adaptive scheme effectively reduces the energy consumption by 20% when compared with the traditional scheme, while guaranteeing the same learning performance.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile devices. Section III develops an energy-efficient LDPC decoding strategy for FL systems by exploiting the relationship between the convergence rate of FL and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, concluding remarks are drawn in Section V.
@@ II System Model and Preliminaries
In this section, we first introduce the wireless FL system. An overview regarding the energy consumption of different components at a mobile unit is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL. For ease of reference, we list the key notations of the system model in Table I.
@@
$\mathbf{w}$ DNN model.
$\Delta\mathbf{w}$ The local model update.
$f$ The loss function of training.
$f^{*}$ The minimum value of the training loss.
${\nabla}f$ The gradient.
$r$ The index of the communication round.
$R$ The total communication rounds.
$e$ The index of the local training step.
$E$ The local training steps.
$\sigma_{L}^{2}$ The variance of local gradients estimator.
$\sigma_{G}^{2}$ The dissimilarity between local and global gradients.
$k$ The index of the mobile device.
$n$ The number of total mobile devices.
$s$ The number of selected clients in every round
$S_{r}$ The set of selected clients in the $r$-th round.
$d$ The dimensions of the DNN model.
$p_{k}$ The data size ratio of the $k$-th client.
$T$ The total training iterations of one client.
$L$ The L-smooth parameter.
$b_{r}$ The bit error rate in $r$-th communication round.
$\eta$ The learning rate.
$M$ The biggest absolute weight value.
$Q$ The LDPC maximum decoding iteration.
(a) Wireless FL system. Multiple mobile clients are coordinated by one remote server through the noisy wireless channels. The communicated parameters are model updates in the uplink and model weights in the downlink.
(b) System operations. Server: aggregation, quantization, LDPC encoding, client selection and broadcasting. Mobile devices: receiving model weights, LDPC decoding,local training and uplink transmission.
Figure 1: Wireless FL in digital communication, system outlook and underlying operations.
@@ II-A Wireless FL
Fig. 1 depicts a wireless FL system, where model training of different clients is coordinated by the server. The FL problem can be formulated as follows [26, 27, 28]:
$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\sum_{k=1}^{n}p_{k}f_{k}(\mathbf{w}),$ (1)
where $\mathbf{w}$ denotes the model parameters, $n$ represents the number of clients, $p_{k}$ is the ratio between the amount of training data samples of the $k$-th client and that of all the clients, and $f_{k}(\mathbf{w})$ is the local loss function of the $k$-th client.
To train a shared global model, the server communicates with the mobile clients for multiple rounds through noisy wireless channels. In each communication round, the global model at the server is broadcast to a set of selected clients. Each of the selected clients then performs local training based on the received global model and transmits the model updates to the server for aggregation. Specifically, in the $r$-th communication round, the global model is quantized by a stochastic quantizer, encoded by a channel code, and broadcast to clients. The model parameters received by the $k$-th client is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{w}_{r}^{\prime},$ (2)
where $\mathbf{w}_{r}^{\prime}$ denotes the noisy version of $\mathbf{w}_{r}$. The impact of BER can be quantified by considering different error bit positions in the model parameters and the detailed analysis is given in Appendix B. A selected client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
$\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi),e=0,\cdots,E-1,k\in\mathcal{S}_{r},$ (3)
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e},\xi)$ denotes the gradient computed from the local mini-batch datasets $\xi$ in the $k$-th client, and $\mathcal{S}_{r}$ denotes the set of selected clients in the $r$-th communication round.
@@
${\Delta\mathbf{w}^{k}_{r}=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0}},$ (4)
which is quantized, encoded, and uploaded to the server. For simplicity, we assume an orthogonal uplink channel is available for each selected client. After receiving these noisy local updates, denoted as $(\Delta\mathbf{w}^{k}_{r})^{\prime}$, the server performs aggregation to update the global model as follows:
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{|\mathcal{S}_{r}|}\sum_{k\in{\mathcal{S}_{r}}}p_{k}(\Delta\mathbf{w}^{k}_{r})^{\prime}.$ (5)
In this work, we focus on the downlink communication rather than the uplink for two reasons. First, our objective is to design channel decoding schemes to save the energy consumption of mobile devices. For the uplink, the channel decoding happens on the server side which falls out of the scope of this study. Second, the server has a continuous power supply and is able to conduct a sufficient number of decoding iterations to achieve satisfactory BER performance.
@@ II-B Energy Consumption Evaluation
In[30], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [31], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [32] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large portion of the overall energy consumption in wireless FL.
@@ II-C LDPC Decoding
Figure 2: Typical LDPC decoder architecture: the variable nodes and check nodes are connected, and the messages are transmitted between them iteratively.
In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. To that end, we first introduce a typical LDPC decoder architecture in Fig. 2, which contains the variable nodes representing the encoded bits to be corrected and check nodes representing the parity check equation. The check nodes exchange information with variable nodes and perform computations iteratively until achieving the maximum decoding iteration [33]. Such iterative processing leads to high energy consumption of LDPC decoding. To save decoding energy, early termination techniques are usually adopted. Specifically, if the parity-check matrix $\mathbf{H}$ and the codeword $\mathbf{x}$ satisfy $\mathbf{Hx}^{\rm{T}}=0$, the decoding process will stop even if the maximum decoding iteration has not been reached[34]. The early termination scheme can save decoding iterations while preserving BER performance. The energy consumption of LDPC decoding depends on the target BER. As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
@@ III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [35], where $T$ is the total number of training epochs. In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
@@
($L$-Smoothness [12, 35]) The local loss function $f_{k}\left(\cdot\right)$ at the $k$-th client is $L$-smooth, i.e., $\forall\mathbf{w},\hat{\mathbf{w}}$, we have $||\nabla f_{k}(\mathbf{w})-\nabla f_{k}(\hat{\mathbf{w}})||\leq L||\mathbf{w}-\hat{\mathbf{w}}||$.
@@
(Unbiased and Bounded-variance Local Gradient Estimator [43]) Let $\xi_{r}^{k}$ be the batch of randomly sampled data at the $k$-th client in the $r$-th communication round. The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{r}^{k})$ is unbiased, i.e., $\mathbb{E}[\nabla f_{k}(\mathbf{w};\xi_{r}^{k})]=\nabla f_{k}(\mathbf{w})$, and has a bounded variance, i.e., $\mathbb{E}[||\nabla f_{k}(\mathbf{w};\xi_{r}^{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
@@
(Bounded Local and Global Gradient Dissimilarity [43]) The global dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
@@
(Unbiased Quantization [12, 28, 35]) The random quantizer Q($\cdot$) is unbiased, i.e., $\mathbb{E}[Q(\mathbf{w})|\mathbf{w}]=\mathbf{w}$.
With Assumptions 1-4, the convergence bound of wireless FL with a given BER is given by the following theorem.

### v2

@@ Abstract
One of the most critical challenges for deploying distributed learning solutions, such as federated learning (FL), in wireless networks is the limited battery capacity of mobile clients. While it is a common belief that the major energy consumption of mobile clients comes from the uplink data transmission, this paper presents a novel finding, namely the channel decoding operation also contributes significantly to the overall energy consumption of mobile clients in FL. Motivated by this new observation, we propose an energy-efficient adaptive channel decoding scheme that leverages the intrinsic robustness of FL to model errors. In particular, the robustness is exploited to reduce the energy consumption of channel decoders at mobile clients by adaptively adjusting the number of decoding iterations. We theoretically prove that wireless FL with communication errors can converge at the same rate as the case with error-free communication as long as the bit error rate (BER) is properly constrained. An adaptive channel decoding scheme is then proposed to improve the energy efficiency of wireless FL systems. Experimental results demonstrate that the proposed method maintains the same learning accuracy while reducing the channel decoding energy consumption by 20% when compared to existing approaches.
@@ I-A Motivation
Deep learning has achieved remarkable success in numerous realms, such as computer vision and natural language processing [1]. However, the conventional centralized learning paradigm encounters the data isolation issue due to the rising concern about privacy [2]. As a result, uploading the distributed data, which may contain privacy-sensitive information, for centralized learning becomes prohibited. To address this issue, federated learning (FL) was proposed [3] to enable the distributed training of deep learning models without sharing the data of users (e.g., IoT devices). However, to fully unlock the potential of FL, we need to tackle several critical challenges, such as data and device heterogeneity [4, 5, 6], as well as the communication bottleneck[7]. Another issue of wireless FL is the energy consumption of mobile clients due to heavy computation and communication workloads for training and transferring the deep neural network (DNN) parameters[8].
@@
Many innovative approaches have been proposed to reduce the energy consumption of mobile clients in FL from different perspectives. For instance, reinforcement learning was utilized in [9] to identify near-optimal client selection and resource allocation to reduce both the convergence time and energy consumption of mobile clients. In [10], the trade-off between learning accuracy and total energy consumption of users was optimized through computational and communication resource allocation. In [11], the transmission and computation energy was minimized by bandwidth allocation and power control. Besides system-level resource management, techniques such as model sparsification [12] and quantization [13, 14, 15, 16, 17, 18] have been proposed to minimize the communication overhead, thus reducing the energy consumption of the transmitters. Authors of [19] minimized the transmission energy consumption of mobile clients while maintaining the learning performance by optimizing the transmit power, user selection and uplink resource allocation. Authors of [20] adopted quantization and periodic aggregation to reduce the transmission volume and energy, which also provided a theoretical framework about the impact of quantization and periodic aggregation on the convergence rate of FL. Model sparsification and quantization techniques are also beneficial for minimizing the processing energy consumption [21, 8]. Authors of [21] proposed to introduce sparsity at training time to reduce the on-device training energy. Authors of [8] reduced the local computation energy by quantization on both DNN weights and activations.
However, most works on energy-efficient FL focused on the energy consumption of mobile transmitters while few of them paid attention to the energy consumption of channel decoders in mobile receivers. As will be shown in Section II-B, channel decoders consume a substantial proportion of energy because of the iterative processes to decode massive DNN parameters in each communication round of wireless FL.
@@
To address the research gap, in this work, we first quantitatively analyze the energy consumption of different components in a mobile unit and show that the channel decoder contributes significantly to the overall energy consumption. Building on this observation, an energy-efficient adaptive channel decoding scheme is proposed for wireless FL systems, by leveraging the fact that FL is robust to model errors due to its statistical learning nature[22]. Furthermore, by extending the convergence analysis in[20] to the scenario of FL with downlink communication erros, we investigate the relationship between the bit error rate (BER) and the convergence rate of wireless FL, and determine the maximum tolerable BER that can maintain the same convergence rate as the case with error-free communication. Finally, we propose a practical approach to adjust the number of channel decoding iterations in each communication round to achieve the maximum tolerable BER, while reducing the energy consumption of mobile clients in wireless FL.
@@ I-B Contributions
This paper investigates the impact of channel decoding on the energy efficiency of wireless FL. Our major contributions are summarized as follows:
  • We theoretically analyze the relationship between BER and the convergence rate of wireless FL. Specially, an upper bound of the average expected gradient norm of the global training loss with respect to BERs over communication rounds is derived by extending the analysis in [20], based on which, the maximum tolerable BER to achieve the same convergence rate as error-free FL is determined.
  • Based on the theoretical analysis, an adaptive low-density parity checking (LDPC) decoding mechanism is proposed, which controls the maximum number of decoding iterations to satisfy the BER requirement in every communication round and maintain the target convergence rate as error-free wireless FL, while reducing the energy consumption of channel decoders in wireless FL.
  • Experimental results on both IID and non-IID datasets show that the proposed adaptive scheme effectively reduces the energy consumption of channel decoders by 20% when compared with the traditional schemes which set a fixed maximum number of channel decoding iterations, while guaranteeing the same learning performance.
@@ I-C Organization
The remainder of this paper is organized as follows. Section II introduces the system model and presents a comprehensive energy consumption analysis for mobile units. Section III develops an energy-efficient adaptive LDPC decoding scheme for wireless FL systems by exploiting the relationship between the convergence rate and the achieved BER of LDPC decoding. Experimental results are presented in Section IV. Finally, conclusions are drawn in Section V.
@@ II System Model and Preliminaries
In this section, we first introduce the wireless FL system. An overview regarding the energy consumption of different components at a mobile unit (i.e., client) is then presented, which highlights the significant contribution of channel decoding to the total energy consumption of wireless FL. For ease of reference, we list the key notations in Table I.
@@
$\mathbf{w}$ DNN model
$\Delta\mathbf{w}_{r}^{k}$ Local model update of client $k$ at the $r$-th communication round
$f$ The loss function of training
$f^{*}$ Minimum value of the global training loss
$R$ The total communication rounds
$E$ Number of local training steps in each communication round
$\sigma_{L}^{2}$ Variance of local gradient estimator
$\sigma_{G}^{2}$ Dissimilarity between local and global gradient
$n$ Number of total clients
$N$ Each DNN parameter is with $N$ bit representation
$d$ Dimension of the DNN model
$T$ Total training iterations, i.e., $T=RE$.
$L$ The $L$-smooth parameter
$b_{r}$ Bit error rate in the $r$-th communication round
$\eta$ Learning rate
$Q$ Maximum number of LDPC decoding iterations
$M_{r}$ Range of the global model parameters in the $r$-th communication round
$\hat{M}$ Maximum range of global model parameters for all communication rounds
(a) Wireless FL system: Multiple mobile clients are coordinated by one remote server through noisy wireless channels, where model updates and model weights are communicated in the uplink and downlink, respectively.
(b) System operations: The server performs model aggregation, digitalization, LDPC encoding, and model broadcasting. Mobile clients perform LDPC decoding, local training, and uploading the model updates.
Figure 1: Wireless FL in digital communication network and its operations.
@@ II-A Wireless FL
Fig. 1(a) depicts a wireless FL system, where model training of different clients is coordinated by a remote server. The FL problem can be formulated as follows [23, 24, 25]:
$\min_{\mathbf{w}\in{\mathbb{R}^{d}}}f(\mathbf{w})\triangleq\frac{1}{n}\sum_{k=1}^{n}f_{k}(\mathbf{w}),$ (1)
where $\mathbf{w}\in{\mathbb{R}^{d}}$ denotes the DNN model with $d$ parameters, $n$ represents the number of mobile clients, $f_{k}\left(\cdot\right)$ is the local loss function of the $k$-th client, and $f\left(\cdot\right)$ is the global loss function.
Figure 2: Digitalization scheme for model parameters. This part gives an example, say [min, max] is digitalized by 3-bits and shows how each value inside is mapped to ”000”, ”001”, ”…”, ”111”.
To train a shared global model, the server communicates with the clients for multiple rounds through noisy wireless channels. As shown in Fig. 1(b), in each communication round, the global model at the server is broadcast to the clients. Each client then performs local training based on the received global model and transmits the model updates after local training to the server for aggregation. Specifically, in the $r$-th communication round, the global model is digitalized by $N$ bits, encoded by a channel code, and broadcast to clients. The model parameter digitalization scheme is illustrated in Fig. 2, where the range of [min, max] is divided into $12^{N}-$ 1 intervals and the width of each interval is $\Delta=\frac{\text{max-min}}{2^{N}-1}$. In this paper, we assume $N$ is large enough so the continuous and digital representations of model parameters have negligible differences. The model parameters received by the $k$-th client in the $r$-th communication round is given as follows:
$\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r},\ k=1,\cdots,n,\ r=0,\cdots,R-1,$ (2)
where $\mathbf{\tilde{w}}_{r}$ denotes the distorted version of $\mathbf{w}_{r}$ due to channel decoding errors. The impact of channel decoding errors can be quantified by considering the different error bit positions in the $N$-bit representations of model parameters, which will be analyzed in Section III-A. Each client updates the received global model by conducting $E$ steps of local mini-batch Stochastic Gradient Descent (SGD) with a learning rate $\eta$ as follows:
$\begin{split}\mathbf{w}^{k}_{r,e+1}&=\mathbf{w}^{k}_{r,e}-\eta{\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k}),\\ &e=0,\cdots,E-1,\ k=1,\cdots,n,\end{split}$ (3)
where ${\nabla}f_{k}(\mathbf{w}^{k}_{r,e};\xi_{r,e}^{k})$ denotes the gradient computed from a mini-batch datasets $\xi_{r,e}^{k}$ at the $k$-th client.
@@
$\begin{split}\Delta\mathbf{w}^{k}_{r}&=\mathbf{w}^{k}_{r,E}-\mathbf{w}^{k}_{r,0},\\ &r=0,\cdots,R-1,\ k=1,\cdots,n,\end{split}$ (4)
which is also digitalized, encoded, and uploaded to the server. For simplicity, we assume an orthogonal uplink channel is available for each client. After receiving these local updates, denoted as $\Delta\mathbf{\tilde{w}}^{k}_{r}$, the server performs model aggregation to update the global model as follows:
$\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum_{k=1}^{n}\Delta\mathbf{\tilde{w}}^{k}_{r},\ r=0,\cdots,R-1.$ (5)
In this work, we focus on the downlink communication rather than the uplink since our objective is to design novel channel decoding schemes to save the energy consumption of mobile clients. For the uplink, the channel decoding happens at the server, where a powerful channel decoder as well as a sufficient number of decoding iterations can always be applied to achieve satisfactory BER performance. As such, we consider $\Delta\mathbf{\tilde{w}}_{r}^{k}=\Delta\mathbf{w}_{r}^{k}$ in the following.
@@ II-B Energy Consumption Evaluation
In[27], a fabricated AI hardware accelerator was reported to achieve the energy efficiency of 13.7 mJ/epoch for 8-bit quantized AlexNet. With five epochs of local training, the processing energy consumption in each communication round is 68.5 mJ. In [28], a fifth-generation (5G) New Radio wireless transmitter and receiver were reported to achieve the energy efficiencies of 47.9 pJ/b and 33.3 pJ/b for transferring uncoded data, respectively, under a communication distance of around 200 meters. Therefore, the energy required to transmit and receive the 8-bit AlexNet with 60 million parameters in a single communication round is about 39 mJ. Assuming a code rate of 1/2, the communication energy consumption for coded transmission of the AlexNet model parameters in each round is 78 mJ. Besides, the LDPC decoder in [29] achieves the energy efficiency of 20.1 pJ/b/iteration, which was measured with 10 decoding iterations on average. Accordingly, the energy consumption of the LDPC decoder for decoding the AlexNet model parameters is 96.5 mJ. The detailed comparison is presented in Table II, which indicates that the energy consumption of LDPC decoding accounts for a large proportion of the overall energy consumption at clients in wireless FL.
@@ II-C LDPC Decoding
In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. In LDPC, the iterative decoding process continues until $Q$ decoding iterations, denoted as the maximum number of decoding iterations, are completed [30]. Such iterative LDPC decoding processing leads to high energy consumption. To save the decoding energy, early termination techniques are usually adopted. Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations has been completed[31]. The early termination scheme can save decoding iterations without compromising the BER performance. The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations. As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
@@ III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [20], where $T$ is the total number of SGD iterations. In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
@@
($L$-Smoothness [13, 20]) The loss function $f\left(\cdot\right)$ is $L$-smooth, i.e., $\forall\mathbf{w},\mathbf{v}$, we have $||\nabla f(\mathbf{w})-\nabla f(\mathbf{v})||\leq L||\mathbf{w}-\mathbf{v}||$.
@@
(Unbiased and Variance-bounded Local Gradient Estimator [39]) Let $\xi_{k}$ denote the batch of randomly sampled data at client $k$. The gradient estimator $\nabla f_{k}(\mathbf{w};\xi_{k})$ is unbiased, i.e., $\mathbb{E}_{\xi_{k}}[\nabla f_{k}(\mathbf{w};\xi_{k})]=\nabla f_{k}(\mathbf{w})$, and has bounded variance, i.e., $\mathbb{E}_{\xi_{k}}[||\nabla f_{k}(\mathbf{w};\xi_{k})-\nabla f_{k}(\mathbf{w})||^{2}]\leq\sigma_{L}^{2}$.
@@
(Bounded Local and Global Gradient Dissimilarity [39]) The dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
@@
(Independent Bit Errors [40]) Each bit in the bitstream of the decoded model parameters encounters error independently.
With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns. However, our empirical results in Fig. 4 show that it is sufficient to focus on cases with at most one bit error of each model parameter. In particular, Fig. 4 compares the actual model error with the analytical model error expression in (7) assuming at most 1-bit error, which match closely with various BER. Therefore, we impose the following assumption to simplify our analysis.
Assumption 5.
(At Most 1-bit Error of each Model Parameter) There is at most 1-bit error in the $N$-bit representation of each model parameter.
Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcasted global model in the following lemma.
Lemma 1.
Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the bias of each dimension in $\mathbf{w}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (6)
The mean square error caused by bit errors is given as
$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{d\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M^{2},$ (7)
where $b$ is the BER.
Proof.
Please refer to Appendix V-A. ∎
(a) BER=$10^{-2}$
(b) BER=$10^{-3}$
(c) BER=$10^{-4}$
Figure 4: Validation of the model error formula by the experiment of vanilla-CNN on Fashion-MNIST dataset. The experimental setting is in Section IV-A. In every communication round, we compute the actual model error between the noisy DNN model and perfect model, and compare with the value from the analytical function in (7). It is shown that the value from analytical function is very close to the actual value, for different BER levels and model range in different communication rounds.
After understanding the effect of BER, we analyze the convergence property of wireless FL with communication errors.
Lemma 2.
With Assumptions 1-5 and the learning rate $\eta$ satisfies
$1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
the average expected gradient norm of $\{\nabla f(\mathbf{\bar{w}}_{r,e})\}$ is bounded as follows:
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n},\end{split}$ (9)
where $f^{*}$ is the minimum global training loss, $\hat{M}$ is the maximum range of DNN weights over all communication rounds, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
Proof.
Please refer to Appendix V-B. ∎
Remark 1.
It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss. The second and third terms indicates how bit errors in each communication round affect learning convergence. The last two terms stem from non-IID data among clients and the random mini-batch sampling.
We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which satisfies (8) with a sufficient large value of $R$.